OpenStudy (anonymous):

Melanie had 2 1/2 bags of sand for a sandbox. The sandbox needs 3 1/4 bags of sand to fill it. Melanie said she still needs 3/4 of a bag to fill it. Is that answer reasonable? A. No, it is not reasonable because 2 1/2 x 3 1/4 = 8 1/8 B. No, it is not reasonable because 2 1/2 + 3 1/4 = 5 3/4 C. Yes, it is reasonable because 3 1/4 - 2 1/2 = 3/4

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

13/4 - 5/2 = 13/4 - 10/4 = 3/4...yep, ur correct

OpenStudy (anonymous):

tx